three electricians can fix 10 short circuits in 2 hours. how much time will four electricians need to fix 15 short circuits?

Found 3 solutions by addingup, ikleyn, MathTherapy:
Answer by addingup(3677)

3 electricians at 2 hours each is 6 man-hours
10 shorts are fixed in 6*60 = 360 man-minutes
360/10 = 36 man-minutes per short
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
For 4 to fix 15:
15*36 = 378 man-minutes
378/4 = 94.5 or 9412 minutes or 1hr3412 minutes

Answer by ikleyn(52778)

.
3 electricians at 2 hours each is 6 man-hours
10 shorts are fixed in 6*60 = 360 man-minutes
360/10 = 36 man-minutes per short
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
For 4 to fix 15:
15*36 = 540 man-minutes
540/4 = 135 minutes, or 2 hours and 15 minutes.


It is typical joint work problem. See the lesson
    - Rate of work problems
in this site.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this textbook under the topic "Rate of work and joint work problems" of the section "Word problems".


Answer by MathTherapy(10552)


three electricians can fix 10 short circuits in 2 hours. how much time will four electricians need to fix 15 short circuits?
Correct answer: